OpenStudy (anonymous):

If a ,b,c are in A.P and a,mb,c are in G.p then a , m^2b,c are in?

OpenStudy (anonymous):

@hartnn @waterineyes

OpenStudy (anonymous):

\[b = \frac{a + c}{2}\] \[mb = \sqrt{ac}\]

OpenStudy (anonymous):

HP ???

OpenStudy (anonymous):

Yup.... Hw

OpenStudy (anonymous):

Yes in HP..

OpenStudy (anonymous):

Hw???

OpenStudy (anonymous):

\[m^2b^2 = ac\] Divide 1 by this..

OpenStudy (anonymous):

\[2b = a + c\] \[m^2b^2 = ac\] Divide this..

OpenStudy (anonymous):

You will get Harmonic Progression..

hartnn (hartnn):

this was the simplest one from u @Yahoo!

OpenStudy (anonymous):

Yup i got it thxxx

OpenStudy (anonymous):

Welcome dear..